OpenStudy (anonymous):

What property is shown below? 6 × 4 = 4 × 6 Associative Property of Multiplication, Commutative Property of Multiplication

Parth (parthkohli):

Commutative property.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow a * b = b * a }\)

OpenStudy (anonymous):

can you please give me all the property's and how they work, like the purple one above, because your smart

Parth (parthkohli):

Commutative property of multiplication.

Parth (parthkohli):

Okay.

Parth (parthkohli):

Distributive property: \(\Large \color{purple}{\rightarrow a(x + y) = ax + ay }\)

Parth (parthkohli):

Associative property: \(\Large \color{purple}{\rightarrow a + ( b + c)= (a + b) + c }\)

Parth (parthkohli):

Commutative property: \(\Large \color{purple}{\rightarrow a + b = b + a }\)

OpenStudy (anonymous):

is that it

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

thank you